OpenStudy (anonymous):

The jurassic zoo charges $5 for adults and $2 for children, the total bill for 144 from a school trip was $396. How many adults and how many children? @Mathematics

OpenStudy (gw2011):

Let A=Adults Let C=Children 5A+2C=396 A+C=144 Multiply this equation by 2 to find A 5A+2C=396 2A+2C=288 Subtract this equation from the above to find A 3A=108 A=36 A+C=144 C=144-36 C=108 Therefore, there are 36 Adults and 108 Children

OpenStudy (anonymous):

thank you so much:)

OpenStudy (gw2011):

You're welcome